Find general solution of the equation
2(1+|cosx|+|cosx|2+ …...............inf) = 2

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
2^{(1+|cosx| + |cosx|^{2}+.....................\infty )} = 2
2^{(\frac{1}{1-|cosx|})} = 2
\frac{1}{1-|cosx|} = 1
1-|cosx| = 1
|cosx| = 0
cosx = 0
x = 2n\pi \pm \frac{\pi }{2}